# Imaging with super-resolution in changing random media

### v1

@@ 1 Introduction
This algorithm extends and improves upon our previous work on imaging through random media using array data. Previously, we addressed imaging through a single unknown random medium, either weakly scattering [Moscoso2024] or strongly scattering [christie2024]. Here we consider the more challenging scenario of imaging through multiple strongly scattering random media, where super-resolution remains achievable as in the single-medium case [christie2024], while introducing essential algorithmic improvements.
@@
Super-resolution refers to the phenomenon where the point spread function of an isolated source in a strongly scattering medium becomes significantly narrower than in a homogeneous medium using the same array imaging system, thereby providing enhanced resolution. While super-resolution in time-reversal has been extensively observed and analyzed [Fink_2000, Borcea_2002], time-reversal differs fundamentally from imaging. Time-reversal is a physical process: recorded array signals are time-reversed and physically re-emitted into the original medium, generating waves that back-propagate and refocus on the sources that produced them.
@@
Time-reversal does not constitute imaging precisely because neither the ambient medium nor the source locations are known, the physical refocusing occurs without forming an image. In strongly scattering media, time-reversal achieves tighter focusing than in homogeneous media (super-resolution) because scattering effectively enlarges the aperture, making the array data appear to originate from a larger, effective array [Borcea_2002], as illustrated schematically in Figure 1.
@@
In contrast, conventional imaging performs back-propagation analytically or numerically, typically assuming a homogeneous ambient medium despite the fact that the recorded data reflect wave propagation through an unknown scattering medium. To achieve super-resolution in imaging, one must therefore estimate the ambient medium properties. This is precisely what we achieved in [christie2024] for a single random medium, developing an algorithm that exploits large, diverse array datasets.
@@
Figure 1: Schematic view of how super-resolution can be explained by the idea of a larger effective aperture [Borcea_2002]. On the left the point spread function in the strongly scattering medium is illustrated and on the right the one obtained using the same array and bandwidth in the homogeneous medium.
@@
The proposed imaging algorithm generalizes directly to the scattering case, that is, when small scattering objects are present in the imaging domain rather than sources. We can do this using the effective source formalism introduced in [Moscoso14].
@@
The challenge in imaging when the ambient random medium is changing is that we do not assume that we know from which random medium the array data originates. We have developed an algorithm that is robust to how different the random media can be, that is, they can range from different, statistically independent random media to ones that are strongly correlated. This requires significant changes to the algorithm used for a single random medium in [christie2024].
@@ 1.1 Array imaging with data from multiple random media
Within each single random medium sensing matrix $\mbox{{${\cal G}$}}^{[i]}$, the columns are relatively incoherent, that is, uncorrelated, reflecting the expected spatial arrangement of the sources in the image window as implied by the expected resolution of the imaging system [borcea2003theory]. Between different random sensing matrices $\mbox{{${\cal G}$}}^{[i]}$ and $\mbox{{${\cal G}$}}^{[j]}$, $i\neq j$, the columns could be arising from significantly correlated media to totally uncorrelated ones. They could, for example, come from slowly changing random media, which are correlated, or from fully independent random media.
@@ 1.2 The four-step imaging algorithm
The second step uses clustering [Ester1996] to distinguish between the estimates of the Green’s function vectors and the extra noise vectors, so that we recover all the columns of $\mbox{{${\cal G}$}}^{B}$ accurately, as described in Section 3.2. However, the columns of $\mbox{{${\cal G}$}}^{B}$ are unordered so they can not be used for imaging because (a) the random medium from which they come is not known, and (b) the points to which they correspond in the image window are also not known.
In a third step, after the denoising by clustering, we separate the columns of $\mbox{{${\cal G}$}}^{B}$ into groups, each of which corresponds to a random medium $i$, $i=1,\dots,L$, using a new graph method as described in Section 3.3. We then use non-metric multidimensional scaling [Yi2003, Oh2010] in each random medium to associate the columns to their corresponding source points in the image window, thus ordering them. This methodology was introduced in imaging in [Moscoso2024] and in [christie2024], and it is described in Section 3.4.
Once the block sensing matrix is determined and then ordered as described, it can be used to produce $\ell_{1}$-norm [chai2014imaging] or $\ell_{2}$-norm (Kirchhoff Migration) [Borcea_2002] images. This is the fourth step.
@@ 2.1 Data generation with the Foldy-Lax model
We use the Foldy-Lax equations [foldy, lax, Ishimaru1978Ch14, martin, let17] to model wave propagation in strongly scattering media. The Foldy-Lax model uses a discrete collection of randomly distributed scatters in an otherwise uniform background to simulate a multiple scattering medium (see Fig. 2).
@@
We note that super-resolution is observed numerically in more complex and realistic multiple scattering models, such as in [let17], where finite size particles are considered rather than point scatterers, using a spherical harmonic expansion and a fast multipole algorithm for solving the resulting system of equations.
@@ 3 Analytical description of the imaging algorithm
In the third step, we cluster the recovered column vectors again into groups $\mbox{{${\cal G}$}}^{[i]}$ corresponding to the random media $i=1,\dots,L$. The clustering procedure is based on the recovered sparse coefficients in $\bm{X}$. This is presented in section 3.3. Finally, in the fourth step in section 3.4, we associate each column vector to a grid point in the image window using non-metric multidimensional scaling. This step is done separately for each, unordered, sensing matrix $\mbox{{${\cal G}$}}^{[i]}$ as in [christie2024].
@@ 3.1.1 Related work
The above problem has been extensively studied over the last decade, but only a few methods offer guarantees for exact recovery, and even these guarantees hold only under highly restrictive models or assumptions on the dictionary and the coefficient matrix. Spielman et al. [Spielman2013], for example, established exact recovery when the dictionary matrix is square and nonsingular and the coefficient matrix has Bernoulli–Gaussian entries. Agarwal et al. [Agarwal2017] proposed an algorithm that guarantees approximate recovery of overcomplete, incoherent dictionaries. With additional postprocessing, arbitrarily accurate estimates of the dictionary can be obtained, provided the sparsity level is of order $O(N^{1/6})$ or lower for large $N$. Novikov et al. [Novikov2023] also developed an algorithm with provable recovery guarantees under a suitable random model for both the dictionary and the coefficient matrices. Their method achieves exact recovery in polynomial time for sparsity levels that are linear in $N$ up to logarithmic factors, representing a significant improvement over previous approaches, including those cited above.
However, for general dictionaries, existing methods lack theoretical guaranties unless a good initialization is available—an important practical limitation. This limitation is also present in the Method of Optimal Directions (MOD) [Engan1999], which converges to the exact solution only when preconditioned with an appropriate initial estimate of the dictionary columns [Agarwal2013]. Otherwise, the algorithm may converge to a suboptimal solution. The key feature of MOD, compared to other dictionary learning methods, is that at each iteration $n$, it updates the dictionary by solving a least-squares problem through the Moore–Penrose pseudoinverse of the current coefficient matrix $\bm{X}_{n}$, provided the corresponding matrices are invertible. MOD does not prescribe a specific method for solving the sparse coding step (14), which is typically handled by algorithms such as Matching Pursuit or Orthogonal Matching Pursuit.
@@ 3.1.2 Introduction of a modified MOD algorithm
    We use one iteration of the Generalized Lagrangian Multiplier Algorithm (GeLMA) [Moscoso_2012], so
@@ 3.2 Improved estimate of sensing matrix by clustering
An improved dictionary is constructed by clustering the columns of the multiple dictionaries found in the previous step using Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [Ester1996] and selecting representatives from each cluster. Suppose that the set to be clustered is $\{\bm{a}_{i}\}_{i=1}^{\ell}$, where $\ell$ is the total number of normalized columns obtained from the previous dictionary learning step. Given a distance function $d$, threshold parameter $\varepsilon$, and minimal number of neighbors parameter $C_{min}$, the DBSCAN algorithm forms and expands clusters about areas with high point density and labels points in low-density areas as noise. If two points are within distance $\varepsilon$, then they are considered neighbors. If a point has $C_{min}$ neighbors, then it is considered a core sample in a cluster. The found columns are clustered as follows.
@@ 3.4 Grid reconstructions using multidimensional scaling
The Green’s function vectors of each $\hat{\mbox{{${\cal G}$}}}^{[i]}$ can be associated with grid points in the image window using the connectivity based sensor localization algorithm in [Yi2003] and analyzed in [Oh2010]. This is a Multi-Dimensional Scaling (MDS) algorithm, with a proxy metric, used in [Moscoso2024] for imaging in a single weakly inhomogeneous random medium and in [christie2024] for imaging in a single scattering medium. To obtain sufficiently stable grid reconstructions in a strongly scattering medium the imaging system must operate with a suitable bandwidth and physical array size, as explained in  [christie2024] but not discussed further here.
The classical MDS algorithm takes as input the pairwise Euclidean distances between points (in our case the Green’s function vectors) and returns the locations of these points in space, up to an overall rotation, translation and scaling. The classical MDS algorithm uses an eigenvalue decomposition [torgerson1952multidimensional]. The location of the points can also be found by minimizing the Frobenius norm between the input distance matrix and that of the pairwise distance of the points to be found using the SMACOF algorithm (”Scaling by MAjorizing a COmplicated Function”) [Kruskal1964, Leeuw1977]. In this paper we use the SMACOF algorithm due to its better performance, but with a proxy metric as follows.
We use the geodesic graph distance of the edge-weighted fixed-radius nearest neighbors graph with radius $R$ as a distance function for MDS. The graph is constructed using the distance function defined in Eq. (18). The cross-correlations in this distance function have the physical interpretation of time reversal experiments [Fink_2000]. If the signals recorded on an array are the (normalized) column vectors of the sensing matrices $\hat{\mbox{{${\cal G}$}}}^{[i]}$, and are re-emitted into the same medium they will focus at the location of the signal source due to the time-reversibility of the wave equation. Thus, computing the cross-correlation with the nearest neighbor is equivalent to using the nearest neighbor for time-reversal. Reconstructing the grid with cross-correlation connectivity information relies on this fundamental property of the wave equation and therefore is very robust to the inhomogeneities of the media.
More specifically, suppose that $\{\hat{\bm{g}}_{i}\}_{i=1}^{{K}}$ is the collection of estimated Green’s vectors to be ordered, that is, associated with grid points in the image window. To reconstruct the grid, we first construct the connectivity graph. To this end, we fix a radius $R$ and define $\hat{\bm{g}}_{i}$ and $\hat{\bm{g}}_{j}$ to be neighbors if $d(\hat{\bm{g}}_{i},\hat{\bm{g}}_{i})<R$. Weight the edge between neighbors $\hat{\bm{g}}_{i}$ and $\hat{\bm{g}}_{j}$ with $d(\hat{\bm{g}}_{i},\hat{\bm{g}}_{j})$. The distance used in the MDS algorithm is the shortest path graph distance which can be found using, for example, Dijkstra’s algorithm [Dijkstra1959].
@@
After running MDS on each individual sensing matrix $\hat{\mbox{{${\cal G}$}}}^{[i]}$ we are left with a reconstructed grid of K{K} points for each realization of the medium $i$, $i=1,\dots,L$. Each reconstructed grid approximates the grid used to generate the data up to a translation, rotation, and scaling. To map the estimated grids onto a target grid we use Procrustes analysis [gower2004]. Specifically, we assume that we know the true location of $4$ anchor points. In numerical simulations we assume that these $4$ anchor points are the corners of the reconstructed grids. For each estimated grid we center and normalize the collection of estimated and known anchor points. Next, we map the standardized $4$ estimate anchor points onto the standardized $4$ known anchor locations using the linear transformation that minimizes the least squares distance between $4$ known anchor points. Finally, we map the full collection of points into the imaging window by performing the Procrustes transformation on the full collection of points.
@@ 3.5 On the coherence of the columns of the block sensing matrix $\mbox{{${\cal G}$}}^{B}\,$
This raises another question: how can sparse dictionary learning succeed when dictionary elements are highly coherent? Our previous work on the role of vicinities in imaging  [MoscosoNovikovPapanicolaouTsogka2020] provides a theoretical framework that addresses this issue. In that work, we showed that imaging can be performed effectively when scatterers are sufficiently separated for each measurement such that their vicinities do not overlap. The same analysis shows that sparse dictionary learning also succeeds under this condition. This requirement is naturally satisfied in changing media: a measurement cannot simultaneously contain signals from sources in two different media configurations.
@@ 4 Numerical simulations
The numerical simulations are conducted in the $C$-band radar regime using the Foldy-Lax model [foldy, lax, let17, martin]. Figure 2 shows a schematic of the imaging system. The imaging system operates with a central frequency of $f_{0}=5$GHz, wavelength $\lambda=6$cm, bandwidth B=1B=1GHz, and a frequency resolution of df=40df=40MHz, resulting in $26$ frequencies. An array composed of $31$ transducers with distance $\lambda$ between them is placed $14$ meters from the center of the image window. The Green’s function vector has dimensions $N=26\cdot 36=806$. The image window dimensions are defined in terms of the central wavelength. The cross-range spacing is $dl_{x}=0.5\lambda$ and the range spacing is $dl_{z}=1.66\lambda$. This produces a total of $K=361$ pixels ($nx=nz=19$). The resolution in a homogeneous medium is estimated at $\lambda L/a\approx 8\lambda$ in cross-range and $5\lambda$ in range. On the other hand, the scattering area between the image window and the array is $10\times 10$ meters and contains $J=400$ point scatterers.

### v2

@@ 1 Introduction
This algorithm extends and improves upon our previous work on imaging through random media using array data. Previously, we addressed imaging through a single unknown random medium, either weakly scattering [22] or strongly scattering [7]. Here we consider the more challenging scenario of imaging through multiple strongly scattering random media, where super-resolution remains achievable as in the single-medium case [7], while introducing essential algorithmic improvements.
@@
Super-resolution refers to the phenomenon where the point spread function of an isolated source in a strongly scattering medium becomes significantly narrower than in a homogeneous medium using the same array imaging system, thereby providing enhanced resolution. While super-resolution in time-reversal has been extensively observed and analyzed [12, 3], time-reversal differs fundamentally from imaging. Time-reversal is a physical process: recorded array signals are time-reversed and physically re-emitted into the original medium, generating waves that back-propagate and refocus on the sources that produced them.
@@
Time-reversal does not constitute imaging precisely because neither the ambient medium nor the source locations are known, the physical refocusing occurs without forming an image. In strongly scattering media, time-reversal achieves tighter focusing than in homogeneous media (super-resolution) because scattering effectively enlarges the aperture, making the array data appear to originate from a larger, effective array [3], as illustrated schematically in Figure 1.
@@
In contrast, conventional imaging performs back-propagation analytically or numerically, typically assuming a homogeneous ambient medium despite the fact that the recorded data reflect wave propagation through an unknown scattering medium. To achieve super-resolution in imaging, one must therefore estimate the ambient medium properties. This is precisely what we achieved in [7] for a single random medium, developing an algorithm that exploits large, diverse array datasets.
@@
Figure 1: Schematic view of how super-resolution can be explained by the idea of a larger effective aperture [3]. On the left the point spread function in the strongly scattering medium is illustrated and on the right the one obtained using the same array and bandwidth in the homogeneous medium.
@@
The proposed imaging algorithm generalizes directly to the scattering case, that is, when small scattering objects are present in the imaging domain rather than sources. We can do this using the effective source formalism introduced in [5].
@@
The challenge in imaging when the ambient random medium is changing is that we do not assume that we know from which random medium the array data originates. We have developed an algorithm that is robust to how different the random media can be, that is, they can range from different, statistically independent random media to ones that are strongly correlated. This requires significant changes to the algorithm used for a single random medium in [7].
@@ 1.1 Array imaging with data from multiple random media
Within each single random medium sensing matrix $\mbox{{${\cal G}$}}^{[i]}$, the columns are relatively incoherent, that is, uncorrelated, reflecting the expected spatial arrangement of the sources in the image window as implied by the expected resolution of the imaging system [4]. Between different random sensing matrices $\mbox{{${\cal G}$}}^{[i]}$ and $\mbox{{${\cal G}$}}^{[j]}$, $i\neq j$, the columns could be arising from significantly correlated media to totally uncorrelated ones. They could, for example, come from slowly changing random media, which are correlated, or from fully independent random media.
@@ 1.2 The four-step imaging algorithm
The second step uses clustering [11] to distinguish between the estimates of the Green’s function vectors and the extra noise vectors, so that we recover all the columns of $\mbox{{${\cal G}$}}^{B}$ accurately, as described in Section 3.2. However, the columns of $\mbox{{${\cal G}$}}^{B}$ are unordered so they can not be used for imaging because (a) the random medium from which they come is not known, and (b) the points to which they correspond in the image window are also not known.
In a third step, after the denoising by clustering, we separate the columns of $\mbox{{${\cal G}$}}^{B}$ into groups, each of which corresponds to a random medium $i$, $i=1,\dots,L$, using a new graph method as described in Section 3.3. We then use non-metric multidimensional scaling [26, 24] in each random medium to associate the columns to their corresponding source points in the image window, thus ordering them. This methodology was introduced in imaging in [22] and in [7], and it is described in Section 3.4.
Once the block sensing matrix is determined and then ordered as described, it can be used to produce $\ell_{1}$-norm [6] or $\ell_{2}$-norm (Kirchhoff Migration) [3] images. This is the fourth step.
@@ 2.1 Data generation with the Foldy-Lax model
We use the Foldy-Lax equations [13, 17, 15, 19, 18] to model wave propagation in strongly scattering media. The Foldy-Lax model uses a discrete collection of randomly distributed scatters in an otherwise uniform background to simulate a multiple scattering medium (see Fig. 2).
@@
We note that super-resolution is observed numerically in more complex and realistic multiple scattering models, such as in [18], where finite size particles are considered rather than point scatterers, using a spherical harmonic expansion and a fast multipole algorithm for solving the resulting system of equations.
@@ 3 Analytical description of the imaging algorithm
In the third step, we cluster the recovered column vectors again into groups $\mbox{{${\cal G}$}}^{[i]}$ corresponding to the random media $i=1,\dots,L$. The clustering procedure is based on the recovered sparse coefficients in $\bm{X}$. This is presented in section 3.3. Finally, in the fourth step in section 3.4, we associate each column vector to a grid point in the image window using non-metric multidimensional scaling. This step is done separately for each, unordered, sensing matrix $\mbox{{${\cal G}$}}^{[i]}$ as in [7].
@@ 3.1.1 Related work
The above problem has been extensively studied over the last decade, but only a few methods offer guarantees for exact recovery, and even these guarantees hold only under highly restrictive models or assumptions on the dictionary and the coefficient matrix. Spielman et al. [27], for example, established exact recovery when the dictionary matrix is square and nonsingular and the coefficient matrix has Bernoulli–Gaussian entries. Agarwal et al. [2] proposed an algorithm that guarantees approximate recovery of overcomplete, incoherent dictionaries. With additional postprocessing, arbitrarily accurate estimates of the dictionary can be obtained, provided the sparsity level is of order $O(N^{1/6})$ or lower for large $N$. Novikov et al. [23] also developed an algorithm with provable recovery guarantees under a suitable random model for both the dictionary and the coefficient matrices. Their method achieves exact recovery in polynomial time for sparsity levels that are linear in $N$ up to logarithmic factors, representing a significant improvement over previous approaches, including those cited above.
However, for general dictionaries, existing methods lack theoretical guaranties unless a good initialization is available—an important practical limitation. This limitation is also present in the Method of Optimal Directions (MOD) [10], which converges to the exact solution only when preconditioned with an appropriate initial estimate of the dictionary columns [1]. Otherwise, the algorithm may converge to a suboptimal solution. The key feature of MOD, compared to other dictionary learning methods, is that at each iteration $n$, it updates the dictionary by solving a least-squares problem through the Moore–Penrose pseudoinverse of the current coefficient matrix $\bm{X}_{n}$, provided the corresponding matrices are invertible. MOD does not prescribe a specific method for solving the sparse coding step (14), which is typically handled by algorithms such as Matching Pursuit or Orthogonal Matching Pursuit.
@@ 3.1.2 Introduction of a modified MOD algorithm
    We use one iteration of the Generalized Lagrangian Multiplier Algorithm (GeLMA) [20], so
@@ 3.2 Improved estimate of sensing matrix by clustering
An improved dictionary is constructed by clustering the columns of the multiple dictionaries found in the previous step using Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [11] and selecting representatives from each cluster. Suppose that the set to be clustered is $\{\bm{a}_{i}\}_{i=1}^{\ell}$, where $\ell$ is the total number of normalized columns obtained from the previous dictionary learning step. Given a distance function $d$, threshold parameter $\varepsilon$, and minimal number of neighbors parameter $C_{min}$, the DBSCAN algorithm forms and expands clusters about areas with high point density and labels points in low-density areas as noise. If two points are within distance $\varepsilon$, then they are considered neighbors. If a point has $C_{min}$ neighbors, then it is considered a core sample in a cluster. The found columns are clustered as follows.
@@ 3.4 Grid reconstructions using multidimensional scaling
The Green’s function vectors of each $\hat{\mbox{{${\cal G}$}}}^{[i]}$ can be associated with grid points in the image window using the connectivity based sensor localization algorithm in [26] and analyzed in [24]. This is a Multi-Dimensional Scaling (MDS) algorithm, with a proxy metric, used in [22] for imaging in a single weakly inhomogeneous random medium and in [7] for imaging in a single scattering medium. To obtain sufficiently stable grid reconstructions in a strongly scattering medium the imaging system must operate with a suitable bandwidth and physical array size, as explained in  [7] but not discussed further here.
The classical MDS algorithm takes as input the pairwise Euclidean distances between points (in our case the Green’s function vectors) and returns the locations of these points in space, up to an overall rotation, translation and scaling. The classical MDS algorithm uses an eigenvalue decomposition [28]. The location of the points can also be found by minimizing the Frobenius norm between the input distance matrix and that of the pairwise distance of the points to be found using the SMACOF algorithm (”Scaling by MAjorizing a COmplicated Function”) [16, 8]. In this paper we use the SMACOF algorithm due to its better performance, but with a proxy metric as follows.
We use the geodesic graph distance of the edge-weighted fixed-radius nearest neighbors graph with radius $R$ as a distance function for MDS. The graph is constructed using the distance function defined in Eq. (18). The cross-correlations in this distance function have the physical interpretation of time reversal experiments [12]. If the signals recorded on an array are the (normalized) column vectors of the sensing matrices $\hat{\mbox{{${\cal G}$}}}^{[i]}$, and are re-emitted into the same medium they will focus at the location of the signal source due to the time-reversibility of the wave equation. Thus, computing the cross-correlation with the nearest neighbor is equivalent to using the nearest neighbor for time-reversal. Reconstructing the grid with cross-correlation connectivity information relies on this fundamental property of the wave equation and therefore is very robust to the inhomogeneities of the media.
More specifically, suppose that $\{\hat{\bm{g}}_{i}\}_{i=1}^{{K}}$ is the collection of estimated Green’s vectors to be ordered, that is, associated with grid points in the image window. To reconstruct the grid, we first construct the connectivity graph. To this end, we fix a radius $R$ and define $\hat{\bm{g}}_{i}$ and $\hat{\bm{g}}_{j}$ to be neighbors if $d(\hat{\bm{g}}_{i},\hat{\bm{g}}_{i})<R$. Weight the edge between neighbors $\hat{\bm{g}}_{i}$ and $\hat{\bm{g}}_{j}$ with $d(\hat{\bm{g}}_{i},\hat{\bm{g}}_{j})$. The distance used in the MDS algorithm is the shortest path graph distance which can be found using, for example, Dijkstra’s algorithm [9].
@@
After running MDS on each individual sensing matrix $\hat{\mbox{{${\cal G}$}}}^{[i]}$ we are left with a reconstructed grid of K{K} points for each realization of the medium $i$, $i=1,\dots,L$. Each reconstructed grid approximates the grid used to generate the data up to a translation, rotation, and scaling. To map the estimated grids onto a target grid we use Procrustes analysis [14]. Specifically, we assume that we know the true location of $4$ anchor points. In numerical simulations we assume that these $4$ anchor points are the corners of the reconstructed grids. For each estimated grid we center and normalize the collection of estimated and known anchor points. Next, we map the standardized $4$ estimate anchor points onto the standardized $4$ known anchor locations using the linear transformation that minimizes the least squares distance between $4$ known anchor points. Finally, we map the full collection of points into the imaging window by performing the Procrustes transformation on the full collection of points.
@@ 3.5 On the coherence of the columns of the block sensing matrix $\mbox{{${\cal G}$}}^{B}\,$
This raises another question: how can sparse dictionary learning succeed when dictionary elements are highly coherent? Our previous work on the role of vicinities in imaging  [21] provides a theoretical framework that addresses this issue. In that work, we showed that imaging can be performed effectively when scatterers are sufficiently separated for each measurement such that their vicinities do not overlap. The same analysis shows that sparse dictionary learning also succeeds under this condition. This requirement is naturally satisfied in changing media: a measurement cannot simultaneously contain signals from sources in two different media configurations.
@@ 4 Numerical simulations
The numerical simulations are conducted in the $C$-band radar regime using the Foldy-Lax model [13, 17, 18, 19]. Figure 2 shows a schematic of the imaging system. The imaging system operates with a central frequency of $f_{0}=5$GHz, wavelength $\lambda=6$cm, bandwidth B=1B=1GHz, and a frequency resolution of df=40df=40MHz, resulting in $26$ frequencies. An array composed of $31$ transducers with distance $\lambda$ between them is placed $14$ meters from the center of the image window. The Green’s function vector has dimensions $N=26\cdot 36=806$. The image window dimensions are defined in terms of the central wavelength. The cross-range spacing is $dl_{x}=0.5\lambda$ and the range spacing is $dl_{z}=1.66\lambda$. This produces a total of $K=361$ pixels ($nx=nz=19$). The resolution in a homogeneous medium is estimated at $\lambda L/a\approx 8\lambda$ in cross-range and $5\lambda$ in range. On the other hand, the scattering area between the image window and the array is $10\times 10$ meters and contains $J=400$ point scatterers.